Kim traveled 2x+50 miles on Monday and 3x-10 miles on Tuesday. Write a polynominal that represents the total distance traveled in 2 days and find the total distance if x=20? Thanks word problems and distance just don't click with me.
Answer by jejess(7)

2x+50 (Monday)
3x-10 (Tuesday)
You need to add monday and tuesday together... therefore the equ. would look like this
Monday+Tuedsy = Total
(2x+50)+(3x-10)= total in two days
x=20 find distance
sub x=20 in where you see x's
(2(20)+50) +(3(20)-10)= total
90+50= 140 miles.
Word problems can be tricky but it is best to write down everything you know first.
Include all your work for full marks on tests, quizes ect.